# Minimax Decision Trees via Martingale Approximations

### v1

@@ 2.3 Non-uniform convergence rates
In other words, under the assumptions in Theorem 5, the MSE has an asymptotic convergence rate of $1/4+\varepsilon$. Unfortunately, we are unable to remove the assumptions for the density of $U$ in Theorem 5 nor give counterexamples. We conjecture that the same conclusion of Theorem 5 holds without those assumptions.
3 The MinimaxSplit algorithm
The goal of this section is to provide a solution to (P-regression) by constructing decision tree algorithms with exponential convergence. Section 3.1 gives the constructions inspired by the minimax martingale construction in Section 2. In traditional regression decision tree settings that minimize (1), models greedily split in order to minimize the sum of variances over the responses of the children nodes (Liu et al., 2023; Breiman et al., 1984), yet we propose to greedily minimize the maximum variance among the children nodes.
In the following, we say that a law on $\mathbb{R}^{d}$ is marginally atomless if its projection onto any of the $d$ dimensions is atomless. We will start from a marginally atomless coupling (that may not arise from empirical measures $\mathbb{P}_{N}$) and prove exponential rates in Section 3.2. Section 3.3 builds upon this case, covering more generally the non-marginally atomless setting and, in particular, the empirical risk bounds. Further numerical analysis will be presented in Section 3.4. All proofs can be found in Appendix A.4.
3.1 Formulation of the algorithms
We begin by recapping the greedy splitting regime in the VarianceSplit algorithm to construct an efficient decision tree. Subsequently, we propose alternative CART algorithms, namely the MinimaxSplit algorithm, along with its variation, the cyclic MinimaxSplit algorithm, and make comparisons against the VarianceSplit algorithm. Roughly speaking, the (cyclic) MinimaxSplit algorithms are analogues of the minimax martingale from Definition 1. Hereafter, we denote by $[d]=\{1,\dots,d\}$ for $d\in\mathbb{N}$. We also need the following notion of a splittable set. If the coupling is defined on a continuous space, the split can always be performed; yet, for atomic measures like empirical measures, we may halt decision splits when there is no probability mass remained. Intuitively, in the following decision tree algorithms, the decision trees only split a node if it corresponds to a splittable set. A similar notion is also introduced by Definition 3.1 of Klusowski and Tian (2024).
Definition 3.1.
Consider a coupling $(\mathbf{X},Y)$ on $\mathbb{R}^{d+1}$. We say a set $A\subseteq\mathbb{R}^{d}$ is $(\mathbf{X},Y)$-non-splittable, or non-splittable, if any of the following occurs:
  • $\mathbb{P}(\mathbf{X}\in A)=0$;
  • $\mathbb{P}(\mathbf{X}\in A)>0$ and $Y\mid\mathbf{X}\in A$ is a constant;
Otherwise, we say that $A$ is $(\mathbf{X},Y)$-splittable, or splittable.
Revisiting the VarianceSplit algorithm.
The greedy VarianceSplit construction (Section 8.4 of Breiman et al. (1984) and Luo and Li (2024)) proceeds by introducing nested partitions $\{\pi_{k}\}_{k\geqslant 0}$ of $\mathbb{R}^{d}$ with axis-aligned borders, in a way that sequentially and greedily minimizes the total risk $\mathbb{E}[(Y-M_{k})^{2}]$ at each step $k$ (while splitting from the fixed partition $\pi_{k-1}$), where $M_{k}$ is taken as the conditional mean of $Y$ given the $\sigma$-algebra generated by the collection ${\mathbbm{1}}_{\{\mathbf{X}\in A\}},\,A\in\pi_{k}$. So, the partitions $\{\pi_{k}\}_{k\geqslant 0}$ of the input space are nested and we require that each $\pi_{k}$ consists of sets that split each element $A\in\pi_{k-1}$ into two hyper-rectangles with axis-aligned borders, unless all covariates and/or all response values on the event $\{\mathbf{X}\in A\}$ are the same, in which case we do not perform split and $A\in\pi_{k}$.
Formally, suppose that a hyper-rectangle $A=[a_{1},b_{1})\times\dots\times[a_{d},b_{d})$ belongs to the partition $\pi_{k-1}$ and is splittable. In particular, $M_{k-1}(\omega)=\mathbb{E}[Y\mid\mathbf{X}\in A]$ if $\mathbf{X}(\omega)\in A$. To find a split of $A$ at depth $k$, the VarianceSplit algorithm looks for a covariate $j\in[d]$ and $x_{j}\in[a_{j},b_{j})$ such that the remaining risk $\mathbb{E}[(Y-M_{k})^{2}{\mathbbm{1}}_{\{\mathbf{X}\in A\}}]$ is the smallest after splitting the $A$ at $x_{j}$ in covariate $j$. In other words, one seeks for
$\begin{split}(j,x_{j})&=\operatorname*{arg\,min}_{\begin{subarray}{c}j\in[d]\\ x_{j}\in[a_{j},b_{j})\end{subarray}}\Big{(}\mathbb{P}(\mathbf{X}\in A,\,X_{j}<x_{j})\mathrm{Var}(Y\mid\mathbf{X}\in A,\,X_{j}<x_{j})\\ &\hskip 56.9055pt+\mathbb{P}(\mathbf{X}\in A,\,X_{j}\geqslant x_{j})\mathrm{Var}(Y\mid\mathbf{X}\in A,\,X_{j}\geqslant x_{j})\Big{)}\\ &=\operatorname*{arg\,min}_{\begin{subarray}{c}j\in[d]\\ x_{j}\in[a_{j},b_{j})\end{subarray}}\Big{(}\mathbb{E}\big{[}(Y-\mathbb{E}[Y\mid\mathbf{X}\in A,\,X_{j}<x_{j}])^{2}{\mathbbm{1}}_{\{\mathbf{X}\in A,\,X_{j}<x_{j}\}}\big{]}\\ &\hskip 56.9055pt+\mathbb{E}\big{[}(Y-\mathbb{E}[Y\mid\mathbf{X}\in A,\,X_{j}\geqslant x_{j}])^{2}{\mathbbm{1}}_{\{\mathbf{X}\in A,\,X_{j}\geqslant x_{j}\}}\big{]}\Big{)}\\ &=:\operatorname*{arg\,min}_{\begin{subarray}{c}j\in[d]\\ x_{j}\in[a_{j},b_{j})\end{subarray}}(V_{\text{left}}+V_{\text{right}}),\end{split}$ (11)
where we break ties arbitrarily. [^3] Strictly speaking, the $\operatorname*{arg\,min}$ in (11) may not always be attained for general couplings $(\mathbf{X},Y)$, unless we assume that $\mathbf{X}$ is marginally either atomless or has a finite support, which will always be the case in our analysis. However, in the general setting, if we allow splits in which atoms can be duplicated and assigned to both nodes (instead of only on the right node as in (11)), $\operatorname*{arg\,min}$ can be attained. In this paper, we will implicitly assume that $\operatorname*{arg\,min}$ is always attained. Consequently, the splittable set $A$ splits into the sets $A_{L}:=[a_{1},b_{1})\times\dots\times[a_{j},x_{j})\times\dots\times[a_{d},b_{d})$ and $A_{R}:=[a_{1},b_{1})\times\dots\times[x_{j},b_{j})\times\dots\times[a_{d},b_{d})$ to form its two descendants in the set $\pi_{k}$, and define
This leads to the desired coupling $(\mathbf{X},Y,\{M_{k}\}_{k\geqslant 0})$, where $\{M_{k}\}_{k\geqslant 0}$ is a martingale. It is also common practice to weigh the terms $V_{\text{left}}$ and $V_{\text{right}}$ in (11) by the sample sizes of the left and right children nodes (resulting in a minimization of $n_{\text{left}}V_{\text{left}}+n_{\text{right}}V_{\text{right}}$), for which we call the Weighted Variance algorithm.
The MinimaxSplit algorithm.
A common feature of the MinimaxSplit and VarianceSplit algorithms is that they both start from a nested sequence of partitions $\{\pi_{k}\}_{k\geqslant 0}$ of $\mathbb{R}^{d}$ consisting of (at-most) binary splits into hyper-rectangles with axis-aligned borders. The $k$-th approximation $M_{k}$ is defined by (13). The two algorithms differ in the way the partition is constructed: in the MinimaxSplit setting, the split is no longer the best split that reduces total variance but instead is the best split that minimizes the maximum variance within the two descendants. Formally, for a splittable set $A=[a_{1},b_{1})\times\dots\times[a_{d},b_{d})\in\pi_{k-1}$, define the split location as
$\begin{split}(j,\hat{x}_{j})&=\operatorname*{arg\,min}_{\begin{subarray}{c}j\in[d]\\ x_{j}\in[a_{j},b_{j})\end{subarray}}\max\Big{\{}\mathbb{E}[(Y-\mathbb{E}[Y\mid\mathbf{X}\in A,\,X_{j}<x_{j}])^{2}{\mathbbm{1}}_{\{\mathbf{X}\in A,\,X_{j}<x_{j}\}}],\\ &\hskip 85.35826pt\mathbb{E}[(Y-\mathbb{E}[Y\mid\mathbf{X}\in A,\,X_{j}\geqslant x_{j}])^{2}{\mathbbm{1}}_{\{\mathbf{X}\in A,\,X_{j}\geqslant x_{j}\}}]\Big{\}}\\ &=\operatorname*{arg\,min}_{\begin{subarray}{c}j\in[d]\\ x_{j}\in[a_{j},b_{j})\end{subarray}}\max\Big{\{}\mathbb{P}(\mathbf{X}\in A,\,X_{j}<x_{j})\mathrm{Var}(Y\mid\mathbf{X}\in A,\,X_{j}<x_{j}),\\ &\hskip 85.35826pt\mathbb{P}(\mathbf{X}\in A,\,X_{j}\geqslant x_{j})\mathrm{Var}(Y\mid\mathbf{X}\in A,\,X_{j}\geqslant x_{j})\Big{\}}\\ &=\operatorname*{arg\,min}_{\begin{subarray}{c}j\in[d]\\ x_{j}\in[a_{j},b_{j})\end{subarray}}\max\Big{\{}V_{\text{left}},V_{\text{right}}\Big{\}},\end{split}$ (14)
where we break ties arbitrarily. Consequently, the splittable set $A$ splits into the two hyper-rectangles $[a_{1},b_{1})\times\dots\times[a_{j},\hat{x}_{j})\times\dots\times[a_{d},b_{d})$ and $[a_{1},b_{1})\times\dots\times[\hat{x}_{j},b_{j})\times\dots\times[a_{d},b_{d})$ to form its two descendants in the set $\pi_{k}$. We call this the MinimaxSplit rule. After obtaining the partition $\pi_{k}$, define $M_{k}$ through (13). This leads to a coupling $(\mathbf{X},Y,\{M_{k}\}_{k\geqslant 0})$, where $\{M_{k}\}_{k\geqslant 0}$ is a martingale.
The cyclic MinimaxSplit algorithm.
The cyclic MinimaxSplit algorithm is a variation of the MinimaxSplit algorithm. Instead of optimizing over all covariates $j\in[d]$ in (14), we cycle through the $d$ covariates as the tree grows. That is, for $k\geqslant 1$, let $j=j_{k}=(k-1\text{ (mod }d))+1\in[d]$. For all splittable sets $A\in\pi_{k-1}$, we split $A$ in the $j$-th coordinate to form its two descendants at depth $k$. The split location is then defined as
where we break ties arbitrarily. Note that if $\mathbf{X}$ is marginally atomless, the minimizer $\hat{x}_{j}$ is essentially unique, in the sense that $\mathbb{P}(\mathbf{X}\in[a_{1},b_{1})\times\dots\times(\hat{x}_{j},\hat{x}_{j}^{\prime})\times\dots\times[a_{d},b_{d}))=0$ between two minimizers $\hat{x}_{j}<\hat{x}_{j}^{\prime}$. Define $M_{k}$ by (13). This leads to the desired coupling $(\mathbf{X},Y,\{M_{k}\}_{k\geqslant 0})$, where again, $\{M_{k}\}_{k\geqslant 0}$ is a martingale aiming to approximate $Y$.
If $d=1$, the cyclic MinimaxSplit algorithm coincides with the MinimaxSplit algorithm. If $d>1$, compared to the MinimaxSplit algorithm, the cyclic MinimaxSplit algorithm is technically more tractable. Moreover, since all features are considered equally, the cyclic MinimaxSplit algorithm appears more robust in recovering the underlying geometry of the regression function in the setting of (4) in low dimensions, which we illustrate numerically in Section 3.4. On the other hand, the cyclic MinimaxSplit algorithm may be less effective in higher dimensions because the major dimensions may not be easily identified or reached in the first few splits.
(At-most-binary) tree representation.
In the context of regression trees, the same development of the tree representation (10) holds, except that the tree may not be a binary tree due to possible non-splittable sets. Instead, there exists a tree $(V,E)$ that is a sub-tree of a binary tree whose vertices are identified as the sets in the nested partitions. With the same notation leading to (10), we have
$\mathbb{E}[(M_{k}-M_{k+1})^{2}]=\sum_{\begin{subarray}{c}v_{k}\in V_{k},\,v_{k+1}\in V_{k+1}\\ v_{k}\sim v_{k+1}\end{subarray}}p_{(v_{k},v_{k+1})}(\ell_{v_{k}}-\ell_{v_{k+1}})^{2}.$
The goal of the next section is to provide a partial answer to (P-general) under the cyclic MinimaxSplit rule. Assuming marginal non-atomicity, we show that up to universal constants and model mis-specification, the approximation error decays exponentially with rate $r^{k}$ for some explicit $r\in(0,1)$ that depends only on dimension $d$.
3.2 Cyclic MinimaxSplit risk bound
We first claim that if $\mathbf{X}$ is marginally atomless and $j\in[d]$, the $L^{2}$ loss
$\mathbb{P}(\mathbf{X}\in A,\,X_{j}<x_{j})\mathrm{Var}(Y\mid\mathbf{X}\in A,\,X_{j}<x_{j})$ (16)
is non-decreasing and continuous in $x_{j}$. To see this, we take $a<b$ and define events $E=E_{1}\cup E_{2}$, where $E_{1}=\{\mathbf{X}\in A,\,X_{j}<a\}$ and $E_{2}=E_{1}^{c}=\{\mathbf{X}\in A,\,a\leqslant X_{j}<b\}.$ By the total variance formula,
Rearranging gives
and therefore (16) (or the first term inside the maximum in (15)) is non-decreasing; and similarly the second term inside the maximum is non-increasing. Moreover, by the marginally atomless property, both quantities are continuous. Therefore, if the $\operatorname*{arg\,min}$ is attained in (15) in the cyclic MinimaxSplit setting, the two terms inside the maximum are equal. Namely,
$\mathbb{P}(\mathbf{X}\in A,\,X_{j}<\hat{x}_{j})\mathrm{Var}(Y\mid\mathbf{X}\in A,\,X_{j}<\hat{x}_{j})=\mathbb{P}(\mathbf{X}\in A,\,X_{j}\geqslant\hat{x}_{j})\mathrm{Var}(Y\mid\mathbf{X}\in A,\,X_{j}\geqslant\hat{x}_{j}).$ (18)
Since the events $A_{L}=\left\{\mathbf{X}\in A,X_{j}<\hat{x}_{j}\right\}$ and $A_{R}=\left\{\mathbf{X}\in A,X_{j}\geqslant\hat{x}_{j}\right\}$ form a partition of $\{\mathbf{X}\in A\}$, we have again by the total variance formula,
$\begin{split}&\hskip 14.22636pt\mathbb{P}(\mathbf{X}\in A,\,X_{j}<\hat{x}_{j})\mathrm{Var}(Y\mid\mathbf{X}\in A,\,X_{j}<\hat{x}_{j})+\mathbb{P}(\mathbf{X}\in A,\,X_{j}\geqslant\hat{x}_{j})\mathrm{Var}(Y\mid\mathbf{X}\in A,\,X_{j}\geqslant\hat{x}_{j})\\ &\leqslant\mathbb{P}(\mathbf{X}\in A)\,\mathbb{E}[(Y-\mathbb{E}[Y\mid\mathbf{X}\in A])^{2}{\mathbbm{1}}_{\{\mathbf{X}\in A\}}].\end{split}$ (19)
Combining (18) and (19), we have
$\begin{split}&\hskip 14.22636pt\max\Big{\{}\mathbb{P}(\mathbf{X}\in A,\,X_{j}<\hat{x}_{j})\,\mathbb{E}[(Y-\mathbb{E}[Y\mid\mathbf{X}\in A,\,X_{j}<\hat{x}_{j}])^{2}{\mathbbm{1}}_{\{\mathbf{X}\in A,\,X_{j}<\hat{x}_{j}\}}],\\ &\hskip 85.35826pt\mathbb{P}(\mathbf{X}\in A,\,X_{j}\geqslant\hat{x}_{j})\,\mathbb{E}[(Y-\mathbb{E}[Y\mid\mathbf{X}\in A,\,X_{j}\geqslant\hat{x}_{j}])^{2}{\mathbbm{1}}_{\{\mathbf{X}\in A,\,X_{j}\geqslant\hat{x}_{j}\}}]\Big{\}}\\ &\leqslant\frac{1}{2}\mathbb{P}(\mathbf{X}\in A)\,\mathbb{E}[(Y-\mathbb{E}[Y\mid\mathbf{X}\in A])^{2}{\mathbbm{1}}_{\{\mathbf{X}\in A\}}].\end{split}$ (20)
In other words, the remaining risk on each descendant is at most half the risk on the parent node. Inductively, we gain a uniform geometric decay of the maximum risk among the nodes at depth $k$:
$\max_{A\in\pi_{k}}\mathbb{P}(\mathbf{X}\in A)\,\mathbb{E}[(Y-\mathbb{E}[Y\mid\mathbf{X}\in A])^{2}{\mathbbm{1}}_{\{\mathbf{X}\in A\}}]\leqslant 2^{-k}\mathrm{Var}(Y).$ (21)
This is the key to controlling the “sizes of the nodes”.
In the case of the VarianceSplit construction, such uniform control is reminiscent—counterexamples exist due to ECP, which explains why assumptions such as the SID condition are favorable to ensure risk decay when using loss functions like (11) (Chi et al., 2022; Mazumder and Wang, 2024). To see the intuition, we consider the event $\{\mathbf{X}\in A\}$ and sub-events $A_{L}=\left\{\mathbf{X}\in A,X_{j}<\hat{x}_{j}\right\}$ and $A_{R}=\left\{\mathbf{X}\in A,X_{j}\geqslant\hat{x}_{j}\right\}$. Applying the total variance formula to this partition yields
$\text{Var}(Y)=$ $\mathbb{E}[\mathrm{Var}(Y\mid\sigma(A_{L}))]+\mathrm{Var}(\mathbb{E}[Y\mid\sigma(A_{L})])$
$\begin{split}=&\text{Var}(Y\mid A_{L})\mathbb{P}(A_{L})+\text{Var}(Y\mid A_{R})\mathbb{P}(A_{R})\\ &+\left(\mathbb{E}[Y\mid A_{L}]-\mathbb{E}[Y\mid A_{R}]\right)^{2}(1-\mathbb{P}(A_{L}))\mathbb{P}(A_{L}).\end{split}$ (22)
From this, we see that when ECP occurs, $\mathbb{P}(A_{L})\approx 1$ or 00, the improvement is dominated by the first two terms of (22): $\text{Var}(Y\mid A_{L})\mathbb{P}(A_{L})+\text{Var}(Y\mid A_{R})\mathbb{P}(A_{R})=V_{\text{left}}+V_{\text{right}}$. Therefore, when ECP occurs, the sum of variances does not necessarily decay fast enough for the VarianceSplit. However, for cyclic MinimaxSplit, $\max\{V_{\text{left}},V_{\text{right}}\}$ always displays geometric decay (21), hence “avoiding” the ECP phenomenon. In other words, the fast decay of the sum of variance is not easily guaranteed from the VarianceSplit construction due to ECP; but the geometric decay of the maximum variance can be assured by the MinimaxSplit construction.
Consider the additive function class
$\mathcal{G}:=\{g(\mathbf{x}):=g_{1}(x_{1})+\dots+g_{d}(x_{d})\},$
where $\mathbf{x}=(x_{1},\dots,x_{d})$. For $g\in\mathcal{G}$, define $\left\lVert g\right\rVert_{\mathrm{TV}}$ as the infimum of $\sum_{i=1}^{d}\left\lVert g_{i}\right\rVert_{\mathrm{TV}}$ over all such additive representations of $g$. (A continuous version of) Theorem 4.2 of Klusowski and Tian (2024) states that if $(\mathbf{X},Y,\{M_{k}\}_{k\geqslant 0})$ is constructed from the VarianceSplit algorithm, then for each $k\geqslant 1$,
In the next result, we show that the $1/k$ rate in (23) can be significantly improved to an exponential rate for the cyclic MinimaxSplit algorithm. This also partially answers (P-general) from the Introduction. Note that the splitting here is binary because of the marginal non-atomicity condition.
Theorem 6.
Suppose that $\mathbf{X}$ is marginally atomless. Let the coupling $(\mathbf{X},Y,\{M_{k}\}_{k\geqslant 0})$ be constructed from the cyclic MinimaxSplit algorithm. Then uniformly for any $\delta>0$ and $k\geqslant 0$,

### v2

@@ A.3 Non-uniform convergence rates
In other words, under the assumptions in Theorem 20, the MSE has an asymptotic convergence rate of $1/4+\varepsilon$”. Unfortunately, we are unable to remove the assumptions for the density of $U$ in Theorem 20 nor give counterexamples. We conjecture that the same conclusion of Theorem 20 holds without those assumptions.
A.4 Proof of Theorem 16
Variance
Lemma A.1.
Let $U$ be an atomless random variable supported on $I=[a,b]$. Then there exists
$u_{*}\in\operatorname*{arg\,min}_{u\in I}\big{(}\mathbb{P}(U\in[a,u))\mathrm{Var}(U\mid U\in[a,u))+\mathbb{P}(U\in[u,b))\mathrm{Var}(U\mid U\in[u,b))\big{)}$
such that
$2u_{*}=\mathbb{E}[U\mid U\leqslant u_{*}]+\mathbb{E}[U\mid U>u_{*}]$ (35)
and moreover, splitting at $u_{*}$ and at $u_{\mathrm{var}}$ defined through (34) have the same effect, [^9] Recall that we break ties in (34) by choosing the largest minimizer. i.e., $\mathbb{P}(U\leqslant u_{*})=\mathbb{P}(U\leqslant u_{\mathrm{var}})$.
Proof.
Let $Q$ be the quantile function of $U$. Since $U$ is atomless, we have if $u=Q(p)$,
$\mathbb{P}(U\in[a,u))\mathrm{Var}(U\mid U\in[a,u))+\mathbb{P}(U\in[u,b))\mathrm{Var}(U\mid U\in[u,b))$
$=p\mathrm{Var}(U\mid U\in[a,Q(p)))+(1-p)\mathrm{Var}(U\mid U\in[Q(p),b)).$
Let $u_{\mathrm{var}}$ be defined through (34) and $p_{\mathrm{var}}=\mathbb{P}(U\leqslant u_{\mathrm{var}})$. Then with the short-hand notation $m_{a}(p)=\mathbb{E}[U\mid U\in[a,Q(p))]$ and $m_{b}(p)=\mathbb{E}[U\mid U\in[Q(p),b)]$, we have
$p_{\mathrm{var}}$ $\in\operatorname*{arg\,min}_{p\in[0,1]}\big{(}p\mathrm{Var}(U\mid U\in[a,Q(p)))+(1-p)\mathrm{Var}(U\mid U\in[Q(p),b))\big{)}$
$=\operatorname*{arg\,max}_{p\in[0,1]}\big{(}p\,m_{a}(p)^{2}+(1-p)m_{b}(p)^{2}\big{)}=:\operatorname*{arg\,max}_{p\in[0,1]}\Psi_{U}(p).$
Observe that for $p\in(0,1)$,
$\frac{\mathrm{d}}{\mathrm{d}p}m_{a}(p)=\frac{Q(p)-m_{a}(p)}{p}\qquad\text{ and }\qquad\frac{\mathrm{d}}{\mathrm{d}p}m_{b}(p)=\frac{m_{b}(p)-Q(p)}{1-p}.$
We then have
$\frac{\mathrm{d}}{\mathrm{d}p}\Psi_{U}(p)$ $=m_{a}(p)^{2}+2m_{a}(p)(Q(p)-m_{a}(p))-m_{b}(p)^{2}+2m_{b}(p)(m_{b}(p)-Q(p))$
$=(m_{b}(p)-m_{a}(p))(m_{a}(p)+m_{b}(p)-2Q(p)).$
Since $U$ is atomless, it is non-degenerate, and hence $m_{b}(p_{\mathrm{var}})-m_{a}(p_{\mathrm{var}})>0$ and $p_{\mathrm{var}}\in(0,1)$. Note that the functions $m_{a},m_{b}$ are differentiable and hence continuous; $Q$ may have jumps but have limits on both sides at every point. Since $p_{\mathrm{var}}\in\operatorname*{arg\,max}_{p\in[0,1]}\Psi_{U}(p)$, there are two cases:
  • $Q$ is continuous at $p_{\mathrm{var}}$. Then $\Psi_{U}^{\prime}(p_{\mathrm{var}})=0$, so $m_{a}(p_{\mathrm{var}})+m_{b}(p_{\mathrm{var}})=2Q(p_{\mathrm{var}})$. In this case, we also have $Q(p_{\mathrm{var}})=u_{\mathrm{var}}$. Therefore, taking $u_{*}=u_{\mathrm{var}}$ gives (35).
  • $Q$ is discontinuous at $p_{\mathrm{var}}$. Then we must have $\Psi_{U}^{\prime}(p_{\mathrm{var}}^{-})\geqslant 0$ and $\Psi_{U}^{\prime}(p_{\mathrm{var}}^{+})\leqslant 0$, where for a generic function $f$ defined on $\mathbb{R}$ we use the notation $f(x^{-})=\lim_{y\to x^{-}}f(y)$ and $f(x^{+})=\lim_{y\to x^{+}}f(y)$. Equivalently,
    $2Q(p_{\mathrm{var}}^{-})\leqslant m_{a}(p_{\mathrm{var}})+m_{b}(p_{\mathrm{var}})\leqslant 2Q(p_{\mathrm{var}}^{+}).$ (36)
    Let $u_{1}=Q(p_{\mathrm{var}}^{-})$ and $u_{2}=Q(p_{\mathrm{var}}^{+})$. Since $Q$ is discontinuous at $p_{\mathrm{var}}$, we have $\mathbb{P}(u_{1}\leqslant U\leqslant u_{2})=0$. Therefore, the map $u\mapsto\mathbb{P}(U\leqslant u)$ is constant on $u_{1},u_{2}]$ and for each $u\in[u_{1},u_{2}]$, $(\mathbb{E}[U\mid U\leqslant u],\mathbb{E}[U\mid U>u])=(m_{a}(p_{\mathrm{var}}),m_{b}(p_{\mathrm{var}}))$. It follows from (36) that (35) holds for some $u_{*}\in[u_{1},u_{2}]$. Using the definition (34) and $p_{\mathrm{var}}=\mathbb{P}(U\leqslant u_{\mathrm{var}})$, we have $u_{\mathrm{var}}=\max_{u\in[u_{1},u_{2}]}u=u_{2}$, which implies that $\mathbb{P}(U\leqslant u_{*})=\mathbb{P}(U\leqslant u_{\mathrm{var}})$.
Therefore, in both cases, the desired claim is verified. ∎